an aeroplane is flying at an altitude of 3000 metres. Two skydivers A and B jump from the plane. A weighs 75 KGS and B weighs 85 kgs. A jumps out without parachute and B jumps with parachute. who will reach the ground faster? that is lesser amount of time why ?explain your answer
1
Expert's answer
2020-04-29T09:45:47-0400

Solution.

"m_A = 75kg;"

"m_B = 85kg;"

"h =3000m;"

"g=9.8m\/s^2;"

f both parachutists A and B were free to fall (without parachute) off the plane, they would move with the acceleration of free fall g.

The time of their motion can be determined from the formula of accelerated motion:

"h = gt^2/2; \implies t = \sqrt{2h/g};"

"t = \sqrt{2\sdot3000m/9.8m/s^2} = 24.74s;"

As you can see, time is independent of mass and it would be the same for parachutist A and B.

As you can see, time is independent of mass and it would be the same for parachutist A and B.

Since, parachutist B flies with a parachute, in addition to gravity, it is subject to the force of air resistance. It is directed vertically upwards and slows down the movement of the parachutist B, thus increasing the time of its fall.

Auger, the sooner the earth will reach parachute A.

Answer: The sooner the parachutis A.